# Goal Inference from Open-Ended Dialog

### v1

@@ IV Experiments
In our experiments, we show that our method enables the robot to perform goal inference to estimate the human’s preference, and then accomplish that task according to it. We conduct experiments for various tasks and human preferences in a text based Grocery Shopping agent domain and an AI2Thor robot simulation domain [27]. We compare our method with two ablation baselines in both of these domains. We also run an isolated inference module experiment on a multiple choice question dataset to see the accuracy of our proposed inference method and to compare the performance between two models of different number of parameters.
IV-A Isolated Inference Intrinsic Evaluation
@@
We check the performance of our proposed goal inference method by testing on a multiple choice dataset [28]. We also compare the performances between using Llama3 8B Instruct and Llama3 70B Instruct models on Hugging Face [29], and check the accuracy with having the “Unspecified Goal” in the goal list. Only open models such as Llama can be used for the inference module because all the logits are available, not just the ones that correspond to the output of text generation.
@@
Each question from the multiple choice dataset has 5 choices. With GPT-4o-mini, we generate four rephrasings for each choice. These 20 choices are the goals in the goal list. For the with “Unspecified Goal” comparison, the “Unspecified Goal” is in the goal list, so there are 21 goals. The sum of the logits that correspond to the words of the original correct choice is taken. The LLM is given the multiple choice question, and each of the 21 goals takes a turn as the “true goal”. If the goal with the highest probability is any of the rephrasings that are derived from the original correct choice, then it is counted as correct. If the most likely goal is any of the rephrasings from the other choices or the “Unspecified Goal”, then it is incorrect.
@@
Figure 3: Isolated Inference Experiments comparing the performance of the Llama3 Instruct 8B vs Llama3 Instruct 70B on only the inference module of the pipeline. Overall, the 8B model has the better performance. The addition of the “Unspecified Goal” does not significantly impact performance.
@@
As seen in Figure 3, we see overall that the 8B Instruct model performs better for both the “Unspecified Goal” not present (94.4%) and “Unspecified Goal” present in the goal list (93.2%) compared to the 70B Instruct model (88.3% and 86.4% respectively). There tiny drop in performance for the “Unspecified Goal” present in the goal list is insignificant, and can be used for determining when to add goals without large impacts on performance. For the remaining experimental results, we use the Meta Llama3 8B Instruct model for inference.
IV-B Experimental Setup
IV-B1 Synthetic Conversation Generation
@@
For our experiments, we simulate human responses with an LLM by conditioning it with a “human profile” at the beginning of the conversation. Both the robot query and the human utterance is generated with GPT4o-mini [30]. All experiments are set at a conversation bound of 20 rounds.
@@
For each round of conversation, the robot generates a question given the robot task description and a transcript of the conversation so far with the human. For the Grocery Shopping agent domain: the task description is “You are a shopping agent that is supposed to make purchases for the human. Your task is to identify a shopping basket that matches the human’s preferences”. For the robot domain: the task description is to “interact with objects in the environment to accomplish the human’s preferences”.
@@
For each round of conversation, given the generated robot query, the human generates a human utterance. We simulate the human with an LLM conditioned on a user human profile consisting of a description of a human’s goal/preference. The human profiles that we do experiments on for the Grocery Shopping domain: five profiles on differing levels of specificity for human profile for gathering ingredients for a chocolate flavored cake (generic chocolate cake, gluten allergy, servings, indecision between flavors, and extra celebratory toppings) along with five other profiles for various meal options are: Italian, SuperBowl, organic, diets for those with anemia. The human profiles for the AI2Thor robot domain experiments are: gathering ingredients for a breakfast sandwich, putting away food, putting away electronic devices, moving valuables into a safe, and gathering cleaning supplies in bathroom.
IV-B2 Inference Module
@@ IV-B3 Action Module
The possible action types for the Grocery Shopping experiment: “search inventory”, “add item to cart”, “remove item from cart”, “buy basket”. The “search inventory” function uses the NLTK package [31] and a Kaggle Grocery Store inventory dataset [32] to implement a simple embedding search by similarity, and retrieves a single most similar item. The shopping basket/cart is represented by a dictionary, so the add item and remove item functions are just dictionary manipulation functions. The “Buy Basket” function is just where the contents of the cart dictionary, total price of the cart are printed out, along with the message “the cart is now purchased”.
@@
For the Grocery Shopping domain, the task is completed and the conversation ends after the “buy basket” action that is called. For the Robot domain, it is after the “task completed” action that is called. The maximum number of conversation rounds is capped at 20. The current state of the cart or successful action history is given to evaluation.
IV-B4 Goal List Evaluations
We check the performance of the Goal Management module through checking if the proposed goal list is reasonable every round, and if the goals that are removed every round are reasonable as well.
IV-B5 Ablation Baselines
@@
IV-B6 LLM Evaluations
We conduct LLM evaluations for our experiments. We evaluate the quality of the generate goal lists. The Goals Reasonable score is assigned out of a overall score of 5, for whether the goals are reasonable given the human utterances, goals information, and high level task for each round of conversation. The Goals Removed Reasonable score is assigned out of a overall score of 3, for whether the goals removed each round are reasonable given the human utterances, goals information, and high level task for each round of conversation.
We also evaluate the outcomes. For the Grocery Shopping Domain, the cart score is assigned out of a score of 3, given the cart, the task of the robot, and the human profile. For the Robot Domain, the successful action transcript for the robot experiment is assigned out of a score of 3, given the final successful action transcript, the task of the robot, and the human profile.
IV-C Results
IV-C1 The Goal List Evaluations
For the Grocery Experiments with our pipeline, the Goals Reasonable score is 97.17% ±$\pm 0.06\%$ and the Goals Removed Reasonable score is 99.09% ±$\pm 0.37\%$. For the Robot Experiments, the Goals Reasonable score is 95.82% ±$\pm 0.48\%$ and the Goals Removed Reasonable score is 93.2% ±$\pm 1.6\%$. For both of these scores, our pipeline does well.
IV-C2 The Outcome Ablation Evaluations
@@
Figure 4: Shopping Domain Ablation Baseline comparisons for LLM evaluations with multiple trials over 10 different human profiles. Our pipeline does better in comparison against a No Goals Baseline that conditions actions on the full dialog. Our pipeline also does better in comparison against a No Inference Baseline that uses a LLM to find the most-likely goal.
@@
Figure 5: Robot Domain Ablation Baseline comparisons for LLM evaluations over multiple trials of five different human profiles. Our pipeline does better in comparison against a No Goals Baseline that conditions actions on the full dialog. Our pipeline also does better in comparison against a No Inference Baseline that uses a LLM to find the most-likely goal.
We compare our method with two ablation baselines for both of our experiment domains. As shown in Figure 4 and Figure 5, our method performs better than both the No Goals Baseline and the No Inference Baseline for both domains. These evaluations are performed by LLM scoring over 5 trials.
V Discussion and Conclusion
We introduced a language-assisted framework for open-ended goal inference that allows for great flexibility in tasks where human preferences are unclear or challenging to specify. We have shown how Language Models can be leveraged to easily edit the support of possible human goals and maintain beliefs over them. We demonstrated that our pipeline can efficiently propose and add new goals based on conversations with the human, and remove goals that are deemed unlikely, undesirable, or unsafe. We also see from our isolated inference evaluation, that overall that the 8B Instruct model performs better compared to the 70B Instruct model.
@@ V Discussion and Conclusion
Future work involves conducting human user studies and see if the LLM evaluations are reflected. While our method shows promise, it currently relies on synthetic conversations generated by LLMs. Future work should validate these results with real human-robot interactions. Other future work can investigate further into less constrained interactions where the human and the agent can take equal roles in providing information, exploration, and inquiry. Our pipeline can also be combined with VLMs, to enable for extracting other representations and amounts of information for both the human and the agent. Our implementation also does not assume that the human needs to answer optimally, and future work can explore cases where the human may be distracted and dividing their attention amongst multiple things, or cases where the human preferences may be slightly altered due to questions asked or persuasion by the agent.
VI Acknowledgements
We thank Jacob Andreas, Phillip Christoffersen, Pinar Ozisik, Belinda Zou Li, Michelle Li, and Andi Peng for helpful conversations. We also thank the reviewers from the HRI-GenAI 2024 Workshop at RSS for their feedback. Rachel Ma was supported by the MIT Presidential Fellowship. Dylan Hadfield-Menell was funded by a Schmidt Sciences AI2050 early career researcher fellowship.

### v2

@@ 4 Experiments
In our experiments, we show that our method enables the robot to perform goal inference to estimate the human’s preference, and then accomplish that task according to it. We conduct experiments for various tasks and human preferences in a text based Grocery Shopping agent domain and an AI2Thor robot simulation domain Kolve et al. (2017). We compare our method with two ablation baselines in both of these domains. We also run an isolated inference module experiment on a multiple choice question dataset to see the accuracy of our proposed inference method and to compare the performance between two models of different number of parameters.
4.1 Isolated Inference Intrinsic Evaluation
@@
We check the performance of our proposed goal inference method by testing on a multiple choice dataset Osmulski (2023). We also compare the performances between using Llama3 8B Instruct and Llama3 70B Instruct models on Hugging Face AI@Meta (2024), and check the accuracy with having the “Unspecified Goal" in the goal list. Only open models such as Llama can be used for the inference module because all the logits are available, not just the ones that correspond to the output of text generation.
@@
Each question from the multiple choice dataset has 5 choices. With GPT-4o-mini, we generate four rephrasings for each choice. These 20 choices are the goals in the goal list. For the with “Unspecified Goal" comparison, the “Unspecified Goal" is in the goal list, so there are 21 goals. The sum of the logits that correspond to the words of the original correct choice is taken. The LLM is given the multiple choice question, and each of the 21 goals takes a turn as the “true goal". If the goal with the highest probability is any of the rephrasings that are derived from the original correct choice, then it is counted as correct. If the most likely goal is any of the rephrasings from the other choices or the “Unspecified Goal", then it is incorrect.
@@
Figure 6: Isolated Inference Experiments comparing the performance of the Llama3 Instruct 8B vs Llama3 Instruct 70B on only the inference module of the pipeline. Overall, the 8B model has the better performance. The addition of the “Unspecified Goal" does not significantly impact performance.
@@
As seen in Figure 6, we see overall that the 8B Instruct model performs better for both the “Unspecified Goal" not present (94.4%) and “Unspecified Goal" present in the goal list (93.2%) compared to the 70B Instruct model (88.3% and 86.4% respectively). There tiny drop in performance for the “Unspecified Goal" present in the goal list is insignificant, and can be used for determining when to add goals without large impacts on performance. For the remaining experimental results, we use the Meta Llama3 8B Instruct model for inference.
4.2 Experimental Setup
Synthetic Conversation Generation:
@@ 4.2 Experimental Setup
For our experiments, we simulate human responses with an LLM by conditioning it with a “human profile” at the beginning of the conversation. Both the robot query and the human utterance is generated with GPT4o-mini 38, with the temperature parameter set to 0. All experiments are set at a conversation bound of 20 rounds.
@@
For each round of conversation, the robot generates a question given the robot task description and a transcript of the conversation so far with the human. For the Grocery Shopping agent domain: the description of the agent and the high level task in the scenario is “You are a shopping agent that is supposed to make purchases for the human. Your task is to identify a shopping basket that matches the human’s preferences”. For the robot domain: the task description is to “interact with objects in the environment to accomplish the human’s preferences”.
@@
For each round of conversation, given the generated robot query, the human generates a human utterance. We simulate the human with an LLM conditioned on a user human profile consisting of a description of a human’s goal/preference. The human profiles that we do experiments on for the Grocery Shopping domain: four profiles on differing levels of specificity for human profile for gathering ingredients for a chocolate flavored cake (generic chocolate cake, gluten allergy, servings, and extra celebratory toppings) along with six other profiles for various meal options are: Italian, SuperBowl, organic, diets for those with anemia, and stocking up a cafe. The human profiles for the AI2Thor robot domain experiments are: gathering ingredients for a breakfast sandwich, putting away food, putting away electronic devices, moving valuables into a safe, and gathering cleaning supplies in bathroom. Since our model does not take in visual input (we only focus on text information from the environment), there is a limitation on what human profiles can be experimented on for the robot domain experiments.
For the check of task completion: for the Grocery Shopper domain, the last sentence of the LLM prompt for the human utterance generation in the Conversation Module, the human is also prompted to “Respond with ‘Buy the basket’ when you deem necessary items that fulfills your human profile and what you have previously said are in the basket after a few rounds of conversations and in their right quantities”. For the Robot domain, the human is prompted with the additional request to “Respond with ‘task completed’ when you deem necessary items that fulfills your human profile”.
Inference Module:
@@
The possible action types for the Grocery Shopping experiment: “search inventory”, “add item to cart”, “remove item from cart”, “buy basket”. The “search inventory” function uses the NLTK package Loper and Bird (2002) and a Kaggle Grocery Store inventory dataset Sakhan (2023) to implement a simple embedding search by similarity, and retrieves a single most similar item. The shopping basket/cart is represented by a dictionary, so the add item and remove item functions are just dictionary manipulation functions. The “Buy Basket" function is just where the contents of the cart dictionary, total price of the cart are printed out, along with the message “the cart is now purchased”.
@@
For the Grocery Shopping domain, the task is completed and the conversation ends after the “buy basket” action that is called. For the Robot domain, it is after the “task completed" action that is called. The maximum number of conversation rounds is capped at 20. The final state of the cart (for the Grocery Shopping domain) or the successful action history (for the robot domain) is passed on to evaluation, either at the end of the 20 rounds or at the point where the task gets completed.
4.3 Ablation Baselines
@@
5 Evaluations
We conduct both LLM evaluations and human evaluations for our experiments. When running our experiments, we have set the temperature for text generation to be 0, and we do a run for each of the human profiles to be evaluated.
For the Grocery Shopping Domain, we evaluate the cart score, which is how relevant/reasonable the items in the final cart is to the human profile and task at hand. For the Robot Domain, we evaluate the successful action score, which is how relevant/reasonable the actions completed in the simulation is to the human profile and the task at hand. Both of these are evaluated based on information given about the task description, the robot description, possible actions that can be enacted in the environment (so that the successful actions list/cart score is understandable), and the human profile. Both of these are scored out of three (where 0 is the least reasonable) and then converted to percentages.
Note of some of the limitations here: we are using Google form surveys for the human evaluators, there is difficulty in how we can present the full information of the dialog transcript cleanly in the form. Since during the experiments, the dialogue is generated from the task description and the human profile, we decided to have the evaluations done without the knowledge of full transcript, and have the human profile as the proxy for it. This and the presence of some other limitations (limited replanning generations in a turn, presence of only certain actions, limited search function, no full knowledge of the grocery inventory), the evaluation prompt for the LLM/human evaluator has the extra request to not penalize for inefficiency and repeats. How we can address these limitations will be expanded upon in the final chapter of the thesis.
5.1 LLM Evaluations
The LLM is prompted with the evaluation questions above, with the additional request to round the score to the nearest .25 and also to provide an explanation for the score. We do 20 trials of LLM scoring for each run (each human profile). For the error bars of the graphs, we rely on calculating pooled standard deviation of the mean.
In occasions when the cart (for the Grocery Shopper Domain) or successful action transcript (for the Robot Domain) is empty (whether before or after purchase), the LLM evaluator will output a hallucinated result. In these cases, we double check the explanation given and the final cart/successful action history. If the explanation acknowledges the empty cart/successful action history or hallucinates one, we make the rating for all of the associated trials for that human profile to be 0.
5.2 Human Evaluations
Human evaluations are conducted through Google form surveys on Prolific Palan and Schitter (2018). Human evaluators are asked questions in the form for both explanations and a rating. Once again, for the error bars of the graphs, we rely on calculating pooled standard deviation of the mean.
With the help of explanations given by the human evaluators, we do not consider evaluations in which we can see that the human evaluators 1) do not understand instructions, 2) do not follow instructions, or 3) spent low effort on the survey (through inserting the same rating or explanation).
5.3 Ablation Results and Discussion
We compare our method with the two ablation baselines for both of our experiment domains as described in Section 4.3.
For the Grocery Shopping text based experiments as shown in Figure 7, our method performs better than the No Goals Baseline and the No Inference Baseline in LLM Evaluations. The human evaluations do not show significant improvement between our pipeline and the No Goals Baseline, but still show a significant improvement over the No Inference baseline. 20 trials of LLM evaluations are collected for each of the human profiles. For the Robot Domain experiments as shown in Figure 8, both the LLM evaluations and Human evaluations do not show significant improvement of our method over the No Goals baseline, but do outperform the No Inference Baseline. For the Grocery Shopper domain, the results are graphed based on 13 human evaluations. For the Robot domain, the results are graphed based on 19 human evaluations.
@@ 5.3 Ablation Results and Discussion
Figure 7: Shopping Domain Ablation Baseline comparisons for LLM evaluations and Human evaluations over 10 different human profiles. Our results show that the inference module improves the performance of the goal editing module as both the LLM and Human evaluations show that our pipeline outperforms the No Inference Baseline significantly. Our comparison with a baseline that conditions actions based on full dialog (No Goals Baseline) that conditions actions on the full dialog shows improvement for the LLM comparisons but no significant improvement in human evaluations.
@@
Figure 8: Robot Domain Ablation Baseline comparisons for LLM evaluations and Human evaluations over five different human profiles. Our pipeline significantly outperforms the No Inference baseline. Our comparison with a No Goals Baseline that conditions actions on the full dialog shows no significant improvement for both human and LLM evaluations.
Our pipeline can do at least as well or better than as if given the whole context of the conversation (the No Goals Baseline). Our pipeline does well with the presence of goals, as they have the capability of summarizing broader ideas that could be relevant to the overall task and human preference, and so our method would be able to allow for performing a number of relevant actions towards the human preferences/task. The difference between the performance of the Grocery Shopping results and the Robot domain results could be an effect of the different types of human profiles that are tested on. The Grocery Shopping domains test a variety of specific profiles for the human. The Robot domain human preferences are limited due to our focus on language information (not directly vision) from the environment.
6 Conclusion
6.1 Future Work
@@
Our pipeline can be adapted to scenarios that allow for even less constrained interactions. The human and the agent can take equal roles in providing information, exploration, and inquiry. More open-ended inquiries from the human or more detailed responses by the robot can be retrieved also with more involved action call designs. Our pipeline can also be combined with VLMs (Vision-Language Models) or other multi-modal methods, to enable for extracting other representations and amounts of information for both the human and the agent. Our implementation also does not assume that the human needs to answer optimally, and future experiments can explore cases where the human may be distracted and dividing their attention between multiple things, or cases where the human preferences may be slightly altered due to questions asked or persuasion by the agent. A better user interface can be used to present experiment data for collecting human evaluations.
Our method currently decides to take an action after each round of conversation between the “human“ and the robot/agent. Ordering of the modules in our pipeline can be experimented with, or experimenting with adding a dialogue action to the actions list for continuing conversation. There can be more flexibility with creating blocks of conversation about particular details associated with human profile or the task at hand, allowing inference calls to be made more efficiently, and we can see if there can be clear improvements of our pipeline over being provided the full conversation context.
6.2 Conclusion
We introduced a language-assisted framework for open-ended goal inference that allows for great flexibility in tasks where human preferences are unclear or challenging to specify. We have shown how Language Models can be leveraged to easily edit the list of possible human goals and maintain beliefs over them. We demonstrated that our pipeline can efficiently propose and add new goals based on conversations with the human, and remove goals that are deemed unlikely, undesirable, or unsafe.